OpenStudy (anonymous):

Solve the system, using substitution. One point for correct steps, one point for correct ordered pair. x = 4y and 2x + 3y = 44

OpenStudy (anonymous):

x=4y 2(4y)+3y=44 8y+3y=44 11y=44 y=44/11=4 y=4 x=4(4) x=16 x=16, y=4

OpenStudy (anonymous):

ahh, thank you so much:)!